OpenStudy (anonymous):

simplified form sqrt 16c^4 - sqrt c^2 + 3 sqrt c^2 + 9c^2 PLease help

OpenStudy (anonymous):

\[\sqrt{16c ^{4}}-\sqrt{c ^{2}}+ 3\sqrt{c ^{2}}+ \sqrt{9c ^{2}}\]

OpenStudy (anonymous):

most of the terms are perfect square, we can extract their square roots and make it all in c variable...

OpenStudy (anonymous):

can you show me?

OpenStudy (anonymous):

... except the 1st term...

OpenStudy (anonymous):

\[\sqrt{16c^4}=\sqrt{16}\times\sqrt{c^4}=4\times c^2\]

OpenStudy (anonymous):

or simply \(4c^2\)

OpenStudy (anonymous):

... we do the same for the rest, you think you can make it?

OpenStudy (anonymous):

lets see.. is 4c^2 = 2c?

OpenStudy (anonymous):

no, we have already removed the radical sign for the 1st term so it's already done, i'm asking for the other terms (2nd to the 4th)...

OpenStudy (anonymous):

the second term would be c..? and the last term would be 9c..? and im ify on the third term ...

OpenStudy (anonymous):

ok for the 2nd term, the last term? 9 is under the radical sign and a perfect square so what is its square root?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

so what is the simplified for the last term?

OpenStudy (anonymous):

3c

OpenStudy (anonymous):

ok you're getting it all right... for the 3rd term, it is c^2 in a radical sign, what is the square root of c^2?

OpenStudy (anonymous):

c

OpenStudy (anonymous):

ok you're correct again, now write all the simplified terms we obtain based on their arrangement above... see if we can further simplify it...

OpenStudy (anonymous):

starting with the 1st term...

OpenStudy (anonymous):

\[4c ^{2}-c+c+3c\]

OpenStudy (anonymous):

you forgot to include the constant 3 in the 3rd term...

OpenStudy (anonymous):

... actually you can simplify the 2nd to the last term since they are all common to c variable...

OpenStudy (anonymous):

\[4c ^{2}-3c+c+3c\]

OpenStudy (anonymous):

like that ?

OpenStudy (anonymous):

it should be \[4c^2-c+3c+3c\]

OpenStudy (anonymous):

so would the final answer be \[4c ^{2}+7c\]

OpenStudy (anonymous):

no, take note the sign of the 2nd term, it is -c

OpenStudy (anonymous):

so \[4c ^{2}+5c\]

OpenStudy (anonymous):

yes you're correct... you got it... :)

OpenStudy (anonymous):

thanks can you help me with one more ?! sorry im so freaking slow..

OpenStudy (anonymous):

ok no problem... :)

OpenStudy (anonymous):

\[5\sqrt{13} +6 -7 \sqrt{13}\]

OpenStudy (anonymous):

i need to find the exact value

OpenStudy (anonymous):

do you think you can simplify further the \(\sqrt{13}~?\)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

if you will look at the expression, do you think you can combine similar terms and perform their operations?

OpenStudy (anonymous):

what are those terms?

OpenStudy (anonymous):

\[13\sqrt{13}\]

OpenStudy (anonymous):

wait no

OpenStudy (anonymous):

\[-2\sqrt{13}\]

OpenStudy (anonymous):

remember you have 6 alone...

OpenStudy (anonymous):

\[-2\sqrt{13} + \sqrt{6}\]

OpenStudy (anonymous):

no it's not, look at the 6, it is under a radical sign?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

then how would you write it now?

OpenStudy (anonymous):

ok it is \[6-2\sqrt{13}.\]